# The hierarchies of identities and closed products for multiple complexes

### v2

@@ 1. Introduction
Ideologically, closed products containing powers of multiple action of mixed differentials represent a geometrical idea of “codimension“ one products. Namely, there are two possibility. First, a product contains a combination of differentials which vanishes due to the critical orders or powers of differentials as the result of application of differentials. Second, an initial candidate to closed products, does not contain maximal orders or powers but after an application of a differential identity, the total maximal powers of elements appear, and a product can be reduced to the first case. The hierarchies of differential identities are of non-trivial nature since we assume that maximal orders of differentials as well as maximal powers of $\mathcal{C}$-elements depend on elements themselves (see explanations in Section 2). Certain non-trivial examples of multiple product identities are given in Sections 35. We provide examples of closed products with respect to multiple products of differentials $d_{a}$ and $d_{\overline{a}}=\overline{d}_{a}$
@@
As for possible applications of the material presented in this paper, we would like to mention computations of higher cohomology for grading-restricted vertex algebras [14], search for more complicated cohomology invariants, and applications in differential geometry and algebraic topology. It would be interesting to study possible applications of invariants we constructed to cohomology of manifolds. In differential geometry there exist various approaches to the construction of cohomology classes (cf., in particular, [17]). We hope to use these techniques to derive counterparts in the cohomology theory of vertex algebras. The results proven in this paper are also useful in computations of cohomology of foliations [4, 6].
@@ 2. Multiple complexes and the main result
[…]
@@
$\ldots\stackrel{{\scriptstyle d^{\ldots,n_{i}-1,\ldots,n_{j},\ldots}_{\ldots,m_{k}+1,\ldots}}}{{\longrightarrow}}C^{\ldots,n_{i},\ldots,n_{j},\ldots}_{\ldots,m_{k},\ldots}\stackrel{{\scriptstyle d^{\ldots,n_{i},\ldots,n_{j},\ldots}_{\ldots,m_{k},\ldots}}}{{\longrightarrow}}C^{\ldots,n_{i}+1,\ldots,n_{j},\ldots}_{\ldots,m_{k}-1,\ldots}\stackrel{{\scriptstyle d^{\ldots,n_{i}+1,\ldots,n_{j},\ldots}_{\ldots,m_{k}-1,\ldots}}}{{\longrightarrow}}\cdots\qquad\quad$
@@
$\ldots\stackrel{{\scriptstyle d^{\ldots,n_{i}-1,\ldots,n_{j}+1,\ldots}_{\ldots,m_{k}+1,\ldots}}}{{\longrightarrow}}C^{\ldots,n_{i},\ldots,n_{j}+1,\ldots}_{\ldots,m_{k},\ldots}\stackrel{{\scriptstyle d^{\ldots,n_{i},\ldots,n_{j}+1,\ldots}_{\ldots,m_{k},\ldots}}}{{\longrightarrow}}C^{\ldots,n_{i}+1,\ldots,n_{j}+1,\ldots}_{\ldots,m_{k}-1,\ldots}\stackrel{{\scriptstyle d^{\ldots,n_{i}+1,\ldots,n_{j}+1,\ldots}_{\ldots,m_{k}-1,\ldots}}}{{\longrightarrow}}\cdots$
@@
[…]
@@
with $\widetilde{\bm{n}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n}}_{i}(\phi_{i})$, $\widetilde{\bm{m}}(\phi)=\sum\limits_{i=1}^{l}{\bm{m}}_{i}(\phi_{i})$ where the summations are component-vise. We set that the result of a $\cdot_{l}$-product which is an element of a $\mathcal{C}$-space, is regular in the domains of definition of all parameters $\Theta^{\bm{n}}_{\bm{m}}$ defining the product. We assume that a $\cdot_{l}$-product is inseparable in pairs of two elements of $\mathcal{C}$ in general. Let us introduce the notations. In the arguments of a multiple product we denote by $(\phi)_{i}$ a $\mathcal{C}$-element placed at the $i$-th position. The notation $\widehat{s}$ means that the value $s$ of the index is omitted. We assume the completeness of a complex $\mathcal{C}$-spaces with respect to the multiple product (2.1). That means that for $\phi$, $\psi\in\mathcal{C}$, placed in the product at the $i$-th and $j$-th positions, from $0=\cdot_{l}(\ldots,(\phi)_{i},\ldots,(\psi)_{j},\ldots)$ it follows that there exists a $\mathcal{C}$-element $\chi$, such that $\psi=\cdot_{l}(\ldots,(\phi)_{i^{\prime}},\ldots,(\chi)_{k},\ldots)$, with some $i^{\prime}$ and $k$. For a $\cdot_{l}$-product we set $\cdot_{l}(\ldots,0,\ldots)=0$. One can also set that $\cdot_{l}(\ldots,{\rm Id}_{\mathcal{C}},\ldots)=\cdot_{l-1}(\ldots,0,\ldots)=0$, where ${\rm Id}_{\mathcal{C}}$ is the identity element in the corresponding $\mathcal{C}$-subspace. In certain cases, a product (2.1) may have coincidences of parameters of multiplied elements of a multiple complex $\mathcal{C}$ spaces. The result of a product may not allow such coincidences leading to possible overcounting of the number of parameters. In order avoid such a possibility, we take into account one coinciding elements/parameters only in the result of a product. For certain $\cdot_{l}$-products of $\mathcal{C}$-elements depending on parameters, (e.g., for a product of vertex operator algebra complexes [14]), it is needed to exclude a number of coinciding arguments and count them only once in the resulting product. Examples of such products can be found in [24, 13] and others. For an element $\phi_{i}\in C^{\bm{n}}_{\bm{m}}$ in a particular $\cdot_{l}$-product let $\bm{r}$ and $\bm{t}$ be the numbers of common parameters corresponding to upper and lower indices for $\phi$ with other elements in a product. In that case, the conditions on indices for a resulting element of a $\cdot_{l}$-product $\phi$ are ${\bm{n}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n}}_{i}(\phi_{i})-{\bm{r}}_{i}(\phi_{i})$, and ${\bm{m}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n}}_{i}(\phi_{i})-{\bm{t}}_{i}(\phi_{i})$. The associativity of a $\cdot_{l}$-product means $(a\cdot_{l}\ldots\cdot_{l^{\prime}}b)\cdot_{l^{\prime\prime}}\ldots\cdot_{l^{\prime\prime\prime}}c=a\cdot_{l}\ldots\cdot_{l^{\prime}}(b\cdot_{l^{\prime\prime}}\ldots\cdot_{l^{\prime\prime\prime}}c)$. for all elements $a$, $b$, $c\in\mathcal{C}$. Note that $l$ can be taken infinite if we assure that a product ×$\times_{i\geq 1}$ is converging with infinite $l$. We call the product formal since elements of a complex $\mathcal{C}$ spaces can be formal (in particular, geometric) objects. Then the result of a $\cdot_{l}$-product is a superposition of formal objects (e.g., Riemann surfaces [24]). In that case, convergence of a product means that the corresponding superposition leads to a well-defined formal object.
@@
In this paper, all the constructions are independent of actual commutation relations for elements for elements inside a $\cdot_{l}$-product. In addition to all that above we assume that in a multiple complex $\mathcal{C}$ spaces there exist a net of subspaces constituting vanishing products of exactly $q$ its elements. That we call an order $q$ ideal $\mathcal{I}(q)\subset\mathcal{C}$. We define a distributed ideal $\mathcal{I}(q)$ of order $q\leq l$ as a union of subsets of $\mathcal{C}$ such that for any set of $q$ elements $\theta_{1}$, $\ldots$, $\theta_{q}\in\mathcal{I}(q)$, distributed in a product, $\cdot_{l}(\ldots$, $\theta_{1}$, $\ldots$, $\theta_{2}$, $\ldots$, $\theta_{q-1}$, …, $\theta_{q}$, )=0$\ldots$)=0… ) = 0. A product (2.1) vanishes if at least one entry of its arguments belonging to $\mathcal{C}$ is zero. Now, let us explain how we understand powers of an element $\phi\in\mathcal{C}$. We denote $(\ldots,(\phi)^{r},\ldots)=(\ldots,(\phi)_{j_{1}},\ldots,(\phi)_{j_{2}},\ldots,(\phi)_{j_{r}},\ldots)$ for $\phi$ placed at the positions $(j_{1},\ldots,j_{r})$ with $r\leq q(\phi)$ where $q(\phi)$ is the maximal power of an element $\phi$, i.e., for $r\geq q(\phi)$, $(\phi)^{r}=0$.
@@
It is taken that a product (2.1) is coherent in indices with respect to all differentials of a complex $\mathcal{C}$. That means that for every set of $\phi_{i}\in C\left(\Theta^{\bm{n_{i}}}_{\bm{m_{i}}}\right)$, $1\leq i\leq l$, ${\bm{n}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n_{i}}}(\phi_{i})$, and ${\bm{m}}(\phi)=\sum\limits_{i=1}^{l}{\bm{m_{i}}}(\phi_{i})$, and the resulting $\phi\in C\left(\Theta^{\bm{n}}_{\bm{m}}\right)$, $\phi=\cdot_{l}(\phi_{1},\ldots,\phi_{l})$, the action of corresponding differential $d$ satisfies Leibniz formula
@@
Note that in general, being applied to the result of a $\cdot_{l}$-product, Leibniz formula represents a map $\mathcal{C}\to\times\mathcal{C}$. Therefore a result of Leibniz formula application vanishes if in each element of the sum (2.2) an element is zero.
@@
we call the union of sets of $\mathcal{C}$-elements $\left\{\Phi_{j}\right\}$, $11\leq j\leq k+$ 1, the completion of a set of elements $\phi_{k}$ in a $\cdot_{l}$-product. Due to the associativity of a $\cdot_{l}$-product we can view each set $\Phi_{j}$ as a single element. The completion (2.3) for a set of elements $\phi_{i}$, $1\leq i\leq k$, with respect to a differential $d_{a}$ is called closed if
@@
where $d_{a,j}$, $j\leq k+1$, acts on $\Phi_{j}$-elements of the completion only. In what follows we will always assume that all completions in $\cdot_{l}$-products are closed with respect to corresponding differentials. Thus, we will skip the completion elements $\Phi_{i},\ldots,\Phi_{k+1}$ of $\mathcal{C}$-elements $\phi_{i}$, $1\leq i\leq k$, so that a $\cdot_{l}$-product will be denoted as $(\ldots,\phi_{1},\ldots,\phi_{k}\ldots)$. Due to the associativity property of the $\cdot_{l}$-product mention above, we will skip the notation $\cdot_{l}$ and denote all completion closed products for various $l$ as (,,)($\ldots$,$\ldots$,$\ldots$)( … , … , … ). Since all actions of $\mathcal{C}$ differentials are coherent with the indexing of $\mathcal{C}$-spaces, we will skip also upper and lower indices from the notations of $d^{n,\kappa}_{m}$.
In general, an arbitrary element $\phi\in\mathcal{C}$ is characterizes by its maximal power $q(\phi)\in\mathbb{Z}_{+}\cup\left\{0\right\}$, i.e., $\left(\ldots,(\phi)_{i_{1}},\ldots,(\phi)_{i_{q(\phi)}},\ldots\right)$ $=0$ with $q$ identical elements $\phi$ distributed in a product (2.1). Next, according to values of maximal orders and powers of a collection of differentials $d_{r_{i}}^{p_{i}}$, $1\leq i\leq k$, a $\mathcal{C}$-element $\phi$ satisfies certain conditions with respect to subsequent actions of differentials $d_{r_{i}}^{p_{i}}$ on $\phi$.
@@
with the corresponding maximal order for the differential $d_{a_{j}}^{r_{j}}$ acting on a $\mathcal{C}$-element $d^{r_{j-1}}_{a_{j-1}}$ldots $d^{r_{1}}_{a_{1}}\phi$, $1\leq j\leq n_{i}$. Note, that due to the property (2.2) and the definition of powers of an element distributed in a $\cdot_{l}$-product, we do not take into account conditions on $\phi\in\mathcal{C}$ of the form $(d_{a_{n}}^{r_{n}}(\ldots(d_{a_{2}}^{r_{2}}(d_{a_{1}}^{r_{1}}\phi)^{q_{1}}))^{q_{2}}\ldots)^{q_{n}}$ since, according to (2.2) it is equivalent to a finite sum of powers of orders of differentials. Note that the definitions of maximal orders and powers above characterize elements of a complex $\mathcal{C}$, and are given in the form of differential/orthogonality relations (6.1) of Section 6. In what follows, we deal with multiple products completion of a number of $\mathcal{C}$-elements closed with respect to corresponding differentials.
@@
Now let us give a proof of Theorem 1.
@@
The idea of the proof is related to geometry. Namely, we show that the identities (2.8) are defined by vanishing products closed with respect a differential $d_{a}$. Geometrically, this corresponds to a separation of “codimension one“ differential forms [15] expressed in terms of multiple products. The main idea to generate differential identities for multiple products of elements of $\mathcal{C}$ is to pick products vanishing due to maximal powers of differentials action. It is easy to see that in order to find a differential identity, one has to pick elements given by multiple products which would give “almost“ the rule (2.7), and would vanish under actions of the differentials $d_{a}$ or $d_{\overline{a}}$. Thus, the set of closed products consists of such “almost“ rule (2.7) elements having “almost“ maximal powers. There are two general ways how to obtain an identity starting from a vanishing product closed under a differential. Let us start with a product containing $k$ entries of $q_{i}(\phi_{i})$-powers of arbitrary $\mathcal{C}$-elements $\phi_{i}$, i.e., $\left(\ldots,(\phi_{i})^{q_{i}},\ldots,(\phi_{j})^{q_{j}},\ldots\right)$, for $0\leq q_{i}<q(\phi_{i})$, with the maximal power $q(\phi_{i})$ of elements $\phi_{i}$ for $1\leq i\neq j\leq k$, and, $q_{j}=q(\phi_{j})$. That product vanishes. Recall that for elements $\phi_{i}$ of $(\phi_{i})^{q_{i}}$ may be distributed as fingle $\phi_{i}$ entries among arguments of the product. Acting by a differential $d_{a}$ according to the formula (2.2) we obtain two types of summands. The action of a differential on the distributed power of $(\phi_{j})^{q(\phi_{j})}$ gives $(d_{a}\phi_{j})$ as one entry, and remaining distributed entries $(\phi_{j})^{q(\phi_{j})-1}$. As we see, since the maximal power $q(\phi_{j})-1$ is dropped now by one, that particular summand does not vanish. There are all together $q(\phi_{j})$ non-vanishing summands of this type. The second type of summands with a differential acting on $(\phi_{i})^{q_{i}}$, $i\neq j$, contains the maximal power of $(\phi_{j})^{q(\phi_{j})}$, and, therefore, vanishes. Another possibility for a vanishing product to be closed is to collect the maximal power of an element at some iteration of differentials application. We thus obtain the first identity of the hierarchy:
@@
This identity relates powers of elements $\phi_{i}$, $1\leq i\leq k$ with the differential $d_{a}\phi_{j}$. To obtain further identities of that branch of the hierarchy, we apply differentials $d_{a}$ or $d_{\overline{a}}$ to (2.9), and take into account (2.5) to derive the next identity in that branch of the hierarchy. Continuing the process, we decrease the powers of elements $(\phi_{i})^{q_{i}}$, and increase powers of differentials $d^{r_{1}}_{a_{1}}\ldots d^{r_{n}}_{a_{n}}\phi_{i}^{q(\phi_{i})}$. Note that by assumption that the maximal order of each differential (i.e., the power when it vanishes) applied depends on the element it acts on. Therefore, the final form of an identity in the hierarchy depends on a sequence of elements $d^{r_{t}}_{a_{t}}(d^{r_{t-1}}_{a_{t-1}}\ldots(d^{r_{1}}_{a_{1}}\phi_{i^{\prime}})))$, $d^{r_{t-1}}_{a_{t-1}}(\ldots(d^{r_{1}}_{a_{1}}\phi_{i^{\prime}}))$, $\ldots$, $d^{r_{1}}_{a_{1}}\phi_{i^{\prime}}$, where $r_{p}(\phi_{p}^{\prime})$ are lower than the maximal orders for corresponding differentials. A sequence of identities stops when at least one order of differentials reaches its maximal value in each summand. The whole hierarchy depends on a set of initial elements $\phi_{i}$. That elements may initially contain orders of differentials of some powers. ∎

### v3

@@ 1. Introduction
Ideologically, closed products containing powers of multiple action of mixed differentials represent a geometrical idea of “codimension“ one products. Namely, there are two possibility. First, a product contains a combination of differentials which vanishes due to the critical orders or powers of differentials as the result of application of differentials. Second, an initial candidate to closed products, does not contain maximal orders or powers but after an application of a differential identity, the total maximal powers of elements appear, and a product can be reduced to the first case. The hierarchies of differential identities are of non-trivial nature since we assume that maximal orders of differentials as well as maximal powers of $\mathcal{C}$-elements depend on elements themselves (see explanations in Section ??). Certain non-trivial examples of multiple product identities are given in Sections ????. We provide examples of closed products with respect to multiple products of differentials $d_{a}$ and $d_{\overline{a}}=\overline{d}_{a}$
@@
As for possible applications of the material presented in this paper, we would like to mention computations of higher cohomology for grading-restricted vertex algebras [15], search for more complicated cohomology invariants, and applications in differential geometry and algebraic topology. It would be interesting to study possible applications of invariants we constructed to cohomology of manifolds. In differential geometry there exist various approaches to the construction of cohomology classes (cf., in particular, [18]). We hope to use these techniques to derive counterparts in the cohomology theory of vertex algebras. The results proven in this paper are also useful in computations of cohomology of foliations [4, 6]. Finally, results of this paper will be useful for computations [26, 27, 28] of topological invariants in quantum field theory via cyclic cohomology methods [7].
@@ 2. Multiple complexes and the main result
$\quad\vdots\qquad\qquad\qquad\qquad\qquad\qquad$$\vdots\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad$
@@
$\ldots\mathrel{{\mathop{\longrightarrow}\limits^{d^{\ldots,n_{i}-1,\ldots,n_{j},\ldots}_{\ldots,m_{k}+1,\ldots}}}}C^{\ldots,n_{i},\ldots,n_{j},\ldots}_{\ldots,m_{k},\ldots}\mathrel{{\mathop{\longrightarrow}\limits^{d^{\ldots,n_{i},\ldots,n_{j},\ldots}_{\ldots,m_{k},\ldots}}}}C^{\ldots,n_{i}+1,\ldots,n_{j},\ldots}_{\ldots,m_{k}-1,\ldots}\mathrel{{\mathop{\longrightarrow}\limits^{d^{\ldots,n_{i}+1,\ldots,n_{j},\ldots}_{\ldots,m_{k}-1,\ldots}}}}\cdots\qquad\quad$
@@
$\ldots\mathrel{{\mathop{\longrightarrow}\limits^{d^{\ldots,n_{i}-1,\ldots,n_{j}+1,\ldots}_{\ldots,m_{k}+1,\ldots}}}}C^{\ldots,n_{i},\ldots,n_{j}+1,\ldots}_{\ldots,m_{k},\ldots}\mathrel{{\mathop{\longrightarrow}\limits^{d^{\ldots,n_{i},\ldots,n_{j}+1,\ldots}_{\ldots,m_{k},\ldots}}}}C^{\ldots,n_{i}+1,\ldots,n_{j}+1,\ldots}_{\ldots,m_{k}-1,\ldots}\mathrel{{\mathop{\longrightarrow}\limits^{d^{\ldots,n_{i}+1,\ldots,n_{j}+1,\ldots}_{\ldots,m_{k}-1,\ldots}}}}\cdots$
@@
$\quad\vdots\qquad\qquad\qquad\qquad\qquad\qquad$$\vdots\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad$
@@
with $\widetilde{\bm{n}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n}}_{i}(\phi_{i})$, $\widetilde{\bm{m}}(\phi)=\sum\limits_{i=1}^{l}{\bm{m}}_{i}(\phi_{i})$ where the summations are component-vise. We set that the result of a $\cdot_{l}$-product which is an element of a $\mathcal{C}$-space, is regular in the domains of definition of all parameters $\Theta^{\bm{n}}_{\bm{m}}$ defining the product. We assume that a $\cdot_{l}$-product is inseparable in pairs of two elements of $\mathcal{C}$ in general. Let us introduce the notations. In the arguments of a multiple product we denote by $(\phi)_{i}$ a $\mathcal{C}$-element placed at the $i$-th position. The notation $\widehat{s}$ means that the value $s$ of the index is omitted. We assume the completeness of a complex $\mathcal{C}$-spaces with respect to the multiple product (LABEL:tugo). That means that for $\phi$, $\psi\in\mathcal{C}$, placed in the product at the $i$-th and $j$-th positions, from $0=\cdot_{l}(\ldots,(\phi)_{i},\ldots,(\psi)_{j},\ldots)$ it follows that there exists a $\mathcal{C}$-element $\chi$, such that $\psi=\cdot_{l}(\ldots,(\phi)_{i^{\prime}},\ldots,(\chi)_{k},\ldots)$, with some $i^{\prime}$ and $k$. For a $\cdot_{l}$-product we set $\cdot_{l}(\ldots,0,\ldots)=0$. One can also set that $\cdot_{l}(\ldots,{\rm Id}_{\mathcal{C}},\ldots)=\cdot_{l-1}(\ldots,0,\ldots)=0$, where ${\rm Id}_{\mathcal{C}}$ is the identity element in the corresponding $\mathcal{C}$-subspace. In certain cases, a product (LABEL:tugo) may have coincidences of parameters of multiplied elements of a multiple complex $\mathcal{C}$ spaces. The result of a product may not allow such coincidences leading to possible overcounting of the number of parameters. In order avoid such a possibility, we take into account one coinciding elements/parameters only in the result of a product. For certain $\cdot_{l}$-products of $\mathcal{C}$-elements depending on parameters, (e.g., for a product of vertex operator algebra complexes [15]), it is needed to exclude a number of coinciding arguments and count them only once in the resulting product. Examples of such products can be found in [25, 14] and others. For an element $\phi_{i}\in C^{\bm{n}}_{\bm{m}}$ in a particular $\cdot_{l}$-product let $\bm{r}$ and $\bm{t}$ be the numbers of common parameters corresponding to upper and lower indices for $\phi$ with other elements in a product. In that case, the conditions on indices for a resulting element of a $\cdot_{l}$-product $\phi$ are ${\bm{n}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n}}_{i}(\phi_{i})-{\bm{r}}_{i}(\phi_{i})$, and ${\bm{m}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n}}_{i}(\phi_{i})-{\bm{t}}_{i}(\phi_{i})$. The associativity of a $\cdot_{l}$-product means $(a\cdot_{l}\ldots\cdot_{l^{\prime}}b)\cdot_{l^{\prime\prime}}\ldots\cdot_{l^{\prime\prime\prime}}c=a\cdot_{l}\ldots\cdot_{l^{\prime}}(b\cdot_{l^{\prime\prime}}\ldots\cdot_{l^{\prime\prime\prime}}c)$. for all elements $a$, $b$, $c\in\mathcal{C}$. Note that $l$ can be taken infinite if we assure that a product $\times_{i\geq 1}$ is converging with infinite $l$. We call the product formal since elements of a complex $\mathcal{C}$ spaces can be formal (in particular, geometric) objects. Then the result of a $\cdot_{l}$-product is a superposition of formal objects (e.g., Riemann surfaces [25]). In that case, convergence of a product means that the corresponding superposition leads to a well-defined formal object.
@@
In this paper, all the constructions are independent of actual commutation relations for elements for elements inside a $\cdot_{l}$-product. In addition to all that above we assume that in a multiple complex $\mathcal{C}$ spaces there exist a net of subspaces constituting vanishing products of exactly $q$ its elements. That we call an order $q$ ideal $\mathcal{I}(q)\subset\mathcal{C}$. We define a distributed ideal $\mathcal{I}(q)$ of order $q\leq l$ as a union of subsets of $\mathcal{C}$ such that for any set of $q$ elements $\theta_{1}$, $\ldots$, $\theta_{q}\in\mathcal{I}(q)$, distributed in a product, $\cdot_{l}(\ldots$, $\theta_{1}$, $\ldots$, $\theta_{2}$, $\ldots$, $\theta_{q-1}$, …, $\theta_{q}$, $\ldots)=0$. A product (LABEL:tugo) vanishes if at least one entry of its arguments belonging to $\mathcal{C}$ is zero. Now, let us explain how we understand powers of an element $\phi\in\mathcal{C}$. We denote $(\ldots,(\phi)^{r},\ldots)=(\ldots,(\phi)_{j_{1}},\ldots,(\phi)_{j_{2}},\ldots,(\phi)_{j_{r}},\ldots)$ for $\phi$ placed at the positions $(j_{1},\ldots,j_{r})$ with $r\leq q(\phi)$ where $q(\phi)$ is the maximal power of an element $\phi$, i.e., for $r\geq q(\phi)$, $(\phi)^{r}=0$.
@@
It is taken that a product (LABEL:tugo) is coherent in indices with respect to all differentials of a complex $\mathcal{C}$. That means that for every set of $\phi_{i}\in C\left(\Theta^{\bm{n_{i}}}_{\bm{m_{i}}}\right)$, $1\leq i\leq l$, ${\bm{n}}(\phi)=\sum\limits_{i=1}^{l}{\bm{n_{i}}}(\phi_{i})$, and ${\bm{m}}(\phi)=\sum\limits_{i=1}^{l}{\bm{m_{i}}}(\phi_{i})$, and the resulting $\phi\in C\left(\Theta^{\bm{n}}_{\bm{m}}\right)$, $\phi=\cdot_{l}(\phi_{1},\ldots,\phi_{l})$, the action of corresponding differential $d$ satisfies Leibniz formula
@@
Note that in general, being applied to the result of a $\cdot_{l}$-product, Leibniz formula represents a map $\mathcal{C}\to\times\mathcal{C}$. Therefore a result of Leibniz formula application vanishes if in each element of the sum (LABEL:leibniz) an element is zero.
@@
we call the union of sets of $\mathcal{C}$-elements $\left\{\Phi_{j}\right\}$, $1\leq j\leq k+1$, the completion of a set of elements $\phi_{k}$ in a $\cdot_{l}$-product. Due to the associativity of a $\cdot_{l}$-product we can view each set $\Phi_{j}$ as a single element. The completion (LABEL:mama) for a set of elements $\phi_{i}$, $1\leq i\leq k$, with respect to a differential $d_{a}$ is called closed if
@@
where $d_{a,j}$, $j\leq k+1$, acts on $\Phi_{j}$-elements of the completion only. In what follows we will always assume that all completions in $\cdot_{l}$-products are closed with respect to corresponding differentials. Thus, we will skip the completion elements $\Phi_{i},\ldots,\Phi_{k+1}$ of $\mathcal{C}$-elements $\phi_{i}$, $1\leq i\leq k$, so that a $\cdot_{l}$-product will be denoted as $(\ldots,\phi_{1},\ldots,\phi_{k}\ldots)$. Due to the associativity property of the $\cdot_{l}$-product mention above, we will skip the notation $\cdot_{l}$ and denote all completion closed products for various $l$ as $(\ldots,\ldots,\ldots)$. Since all actions of $\mathcal{C}$ differentials are coherent with the indexing of $\mathcal{C}$-spaces, we will skip also upper and lower indices from the notations of $d^{n,\kappa}_{m}$.
In general, an arbitrary element $\phi\in\mathcal{C}$ is characterizes by its maximal power $q(\phi)\in\mathbb{Z}_{+}\cup\left\{0\right\}$, i.e., $\left(\ldots,(\phi)_{i_{1}},\ldots,(\phi)_{i_{q(\phi)}},\ldots\right)$ =0=0 with $q$ identical elements $\phi$ distributed in a product (LABEL:tugo). Next, according to values of maximal orders and powers of a collection of differentials $d_{r_{i}}^{p_{i}}$, $1\leq i\leq k$, a $\mathcal{C}$-element $\phi$ satisfies certain conditions with respect to subsequent actions of differentials $d_{r_{i}}^{p_{i}}$ on $\phi$.
@@
with the corresponding maximal order for the differential $d_{a_{j}}^{r_{j}}$ acting on a $\mathcal{C}$-element $d^{r_{j-1}}_{a_{j-1}}$ $\ldots$ $d^{r_{1}}_{a_{1}}\phi$, $1\leq j\leq n_{i}$. Note, that due to the property (LABEL:leibniz) and the definition of powers of an element distributed in a $\cdot_{l}$-product, we do not take into account conditions on $\phi\in\mathcal{C}$ of the form $(d_{a_{n}}^{r_{n}}(\ldots(d_{a_{2}}^{r_{2}}(d_{a_{1}}^{r_{1}}\phi)^{q_{1}}))^{q_{2}}\ldots)^{q_{n}}$ since, according to (LABEL:leibniz) it is equivalent to a finite sum of powers of orders of differentials. Note that the definitions of maximal orders and powers above characterize elements of a complex $\mathcal{C}$, and are given in the form of differential/orthogonality relations (LABEL:groby) of Section ??. In what follows, we deal with multiple products completion of a number of $\mathcal{C}$-elements closed with respect to corresponding differentials.
@@
Now let us give a proof of Theorem ??.
@@
The idea of the proof is related to geometry. Namely, we show that the identities (2.8) are defined by vanishing products closed with respect a differential $d_{a}$. Geometrically, this corresponds to a separation of “codimension one“ differential forms [16] expressed in terms of multiple products. The main idea to generate differential identities for multiple products of elements of $\mathcal{C}$ is to pick products vanishing due to maximal powers of differentials action. It is easy to see that in order to find a differential identity, one has to pick elements given by multiple products which would give “almost“ the rule (LABEL:pirdo), and would vanish under actions of the differentials $d_{a}$ or $d_{\overline{a}}$. Thus, the set of closed products consists of such “almost“ rule (LABEL:pirdo) elements having “almost“ maximal powers. There are two general ways how to obtain an identity starting from a vanishing product closed under a differential. Let us start with a product containing $k$ entries of $q_{i}(\phi_{i})$-powers of arbitrary $\mathcal{C}$-elements $\phi_{i}$, i.e., $\left(\ldots,(\phi_{i})^{q_{i}},\ldots,(\phi_{j})^{q_{j}},\ldots\right)$, for $0\leq q_{i}<q(\phi_{i})$, with the maximal power $q(\phi_{i})$ of elements $\phi_{i}$ for $1\leq i\neq j\leq k$, and, $q_{j}=q(\phi_{j})$. That product vanishes. Recall that for elements $\phi_{i}$ of $(\phi_{i})^{q_{i}}$ may be distributed as fingle $\phi_{i}$ entries among arguments of the product. Acting by a differential $d_{a}$ according to the formula (LABEL:leibniz) we obtain two types of summands. The action of a differential on the distributed power of $(\phi_{j})^{q(\phi_{j})}$ gives $(d_{a}\phi_{j})$ as one entry, and remaining distributed entries $(\phi_{j})^{q(\phi_{j})-1}$. As we see, since the maximal power $q(\phi_{j})-1$ is dropped now by one, that particular summand does not vanish. There are all together $q(\phi_{j})$ non-vanishing summands of this type. The second type of summands with a differential acting on $(\phi_{i})^{q_{i}}$, $i\neq j$, contains the maximal power of $(\phi_{j})^{q(\phi_{j})}$, and, therefore, vanishes. Another possibility for a vanishing product to be closed is to collect the maximal power of an element at some iteration of differentials application. We thus obtain the first identity of the hierarchy:
@@
This identity relates powers of elements $\phi_{i}$, $1\leq i\leq k$ with the differential $d_{a}\phi_{j}$. To obtain further identities of that branch of the hierarchy, we apply differentials $d_{a}$ or $d_{\overline{a}}$ to (LABEL:pordo), and take into account (2.5) to derive the next identity in that branch of the hierarchy. Continuing the process, we decrease the powers of elements $(\phi_{i})^{q_{i}}$, and increase powers of differentials $d^{r_{1}}_{a_{1}}\ldots d^{r_{n}}_{a_{n}}\phi_{i}^{q(\phi_{i})}$. Note that by assumption that the maximal order of each differential (i.e., the power when it vanishes) applied depends on the element it acts on. Therefore, the final form of an identity in the hierarchy depends on a sequence of elements $d^{r_{t}}_{a_{t}}(d^{r_{t-1}}_{a_{t-1}}\ldots(d^{r_{1}}_{a_{1}}\phi_{i^{\prime}})))$, $d^{r_{t-1}}_{a_{t-1}}(\ldots(d^{r_{1}}_{a_{1}}\phi_{i^{\prime}}))$, $\ldots$, $d^{r_{1}}_{a_{1}}\phi_{i^{\prime}}$, where $r_{p}(\phi_{p}^{\prime})$ are lower than the maximal orders for corresponding differentials. A sequence of identities stops when at least one order of differentials reaches its maximal value in each summand. The whole hierarchy depends on a set of initial elements $\phi_{i}$. That elements may initially contain orders of differentials of some powers. ∎